# Causal Classification of Pathological Misner-Type Spacetimes

### v3

@@ 8. Conclusion
A central result of this work is the formal proof of the isocausality of the Misner-type family at the covering level. In Proposition 7.1 we showed that the universal covers of the Misner, pseudo-Schwarzschild, and pseudo-Reissner-Nordström spacetimes are pairwise isocausal via explicit smooth bijections whose differentials preserve and reflect all future-directed causal cones. This isocausality also holds on suitable open, causally regular regions of the compactified spacetimes where the chronology-violating sets and Cauchy horizons are excised. Furthermore, we identified an explicit and checkable deck-equivariance criterion under which the isocausality descends to the full compactified spacetimes. When this criterion fails, Corollary 1 and Remark 7.2 show that the covering-space result does not imply global isocausality, and one should expect only a one-way causal relation between the compactified models. This distinction between local/covering and global/quotient behavior is essential for understanding the true strength of the causal equivalences in these examples.
@@ Acknowledgement.
Parts of this work were completed during my time as a member of Richard Schoen’s research group at the University of California, Irvine. However, the research presented here originated from a project conducted under the supervision of Kip S. Thorne during 2016-2018. I am profoundly grateful to Kip S. Thorne for his invaluable advice, guidance, supportive remarks, and encouragement throughout the course of this work. I would also like to express my deep appreciation for his inspiring spirit of intellectual adventure in research and the insightful discussions, particularly on Misner space, which have been truly indispensable to this work. Furthermore, I am very grateful to the referees for their thorough and constructive feedback.

### v4

@@ 8. Conclusion
A central result of this work is the formal proof of the isocausality of the Misner-type family at the covering level. In Proposition 7.1 we showed that the universal covers of the Misner, pseudo-Schwarzschild, and pseudo-Reissner-Nordström spacetimes are pairwise isocausal via explicit smooth bijections whose differentials preserve future-directed causal cones. This isocausality also holds on suitable open, causally regular regions of the compactified spacetimes where the chronology-violating sets and Cauchy horizons are excised. Furthermore, we identified an explicit and checkable deck-equivariance criterion under which the isocausality descends to the full compactified spacetimes. When the equivariance degree satisfies $|k|=1$, global isocausality follows. If instead $|k|>1$, or if equivariance fails altogether, then Corollary 1 and Remark 7.2 show that one obtains only a one-way causal relation between the compactified models. This distinction between local/covering and global/quotient behavior is essential for understanding the true strength of the causal equivalences in these examples.
@@ Acknowledgement.
Parts of this work were completed during my time as a member of Richard Schoen’s research group at the University of California, Irvine, and at the Simons Center for Geometry and Physics, Stony Brook University. However, the research presented here originated from a project conducted under the supervision of Kip S. Thorne during 2016-2018. I am profoundly grateful to Kip S. Thorne for his invaluable advice, guidance, supportive remarks, and encouragement throughout the course of this work. I would also like to express my deep appreciation for his inspiring spirit of intellectual adventure in research and the insightful discussions, particularly on Misner space, which have been truly indispensable to this work. Furthermore, I am very grateful to the referees for their thorough and constructive feedback.